in a poker hand consisting of 5 cards, find the probability
of holding 2 aces and 3 jacks?

Answer by Edwin McCravy(20054)

4 Aces CHOOSE 2 = 4C2 = (4∙3)/(2∙1) = 12/2 = 6 ways

times:

4 Jacks CHOOSE 3 = 4C3 = (4∙3∙2)/(3∙2∙1) = 24/6 = 4 ways

That's 6×4 = 24 successful ways

52 Cards CHOOSE 5 = (52∙51∙50∙49∙48)/(5∙4∙3∙2∙1) = 311875200/120 =

2598960 possible ways.

Probability = 24 successful ways out of 2598960 possible ways = 
24/2598960 = 1/108290

Edwin